OpenStudy (anonymous):

If f(x) = x2 - 12x + 35, solve for x in the equation f(x) = 3.

OpenStudy (s3a):

3 = x^2 - 12x + 35 0 = x^2 - 12x + 35 - 3 0 = x^2 - 12x + 32 quadratic formula: x = [-b +/- sqrt(b^2 - 4ac)]/(2a) where a = 1, b = -12, c = 32